OpenStudy (hybrik):

Find x+y+z if x(square root of y)=108 y(square root of z)=80(square root of 3) z(square root of x)=225(square root of 3)

OpenStudy (hybrik):

in other terms: \[x \sqrt{y}=108, y \sqrt{z}=80\sqrt{3}, and, z \sqrt{x}=225\sqrt{3}\]

OpenStudy (hybrik):

@ganeshie8

OpenStudy (hybrik):

@Hero

OpenStudy (hybrik):

@welshfella

OpenStudy (hybrik):

any ideas?

OpenStudy (hybrik):

@undeadknight26

undeadknight26 (undeadknight26):

This is mind boggling o.o

OpenStudy (hybrik):

I got 6 minutes then i have lunch.

undeadknight26 (undeadknight26):

Getting confused from just looking at it.

OpenStudy (hybrik):

its a algebra 2 question i believe

undeadknight26 (undeadknight26):

You pretty much have to do them all at once...And the x(sqrt y) thing confuses me.

OpenStudy (hybrik):

i got lunch, i was thinking to square the whole equations

OpenStudy (hybrik):

any ideas?

OpenStudy (freckles):

\[x \sqrt{y}=108 \\ y \sqrt{z}=80 \sqrt{3} \\ z \sqrt{x}=225 \sqrt{3}\] this is right?

OpenStudy (hybrik):

ye

OpenStudy (freckles):

and is x and y and z integers are real?

OpenStudy (hybrik):

I don't know :|

OpenStudy (freckles):

and is x and y and z integers or real? *

OpenStudy (hybrik):

I just want to learn how to do these questions, because I like to learn.

OpenStudy (freckles):

I understand that but I was asking if you knew anything about the variables

OpenStudy (freckles):

I guess we will assume they are real

OpenStudy (hybrik):

ok

OpenStudy (freckles):

\[x^2 y=108^2 \\ y^2 z=80^2(3) \\ z^2 x=225^2 (3) \\ \text{ here I just squared both sides of equation equation } \] \[y=\frac{108^2}{x^2} \text{ solved first equation for } y \\ \text{ replacing } y \text{ \in second equation with this } \\ (\frac{108^2}{x^2})^2z=80^2(3) \\ \text{ by law of exponents you have } \frac{108^4}{x^4} z=80^2 (3) \\ \text{ multiplying both sides by } \frac{x^4}{108^4} \\ z=\frac{80^2(3)x^4}{108^4} \\ \text{ so replacing the } z \text{ \in the last equation with this } \\ (\frac{80^2(3)x^4}{108^4})^2x=225^2(3) \text{ you can solve this one for } x \text{ now }\] this seems like a pretty long way there might be a shorter way but you can solve the last equation for x then go back and find z and then go back and find y then add them up to find x+y+z

OpenStudy (hybrik):

This is taking a bit of time, need 2 minutes.

OpenStudy (freckles):

@ganeshie8 do you see any kind of short cut here

OpenStudy (hybrik):

x=27

OpenStudy (hybrik):

y=16

OpenStudy (freckles):

cool I got x=27 too

OpenStudy (hybrik):

Can't find Z though..

OpenStudy (hybrik):

5 radical 3

OpenStudy (freckles):

\[z=\frac{80^2 (3)x^4}{108^4}\]

OpenStudy (hybrik):

x+y+z = 42+\[5\sqrt{3}\]

OpenStudy (freckles):

\[y=\frac{108^2}{x^2} \\ y=\frac{108^2}{(27)^2}=16 \\ z=\frac{80^2(3)x^4}{108^4} \\ z=\frac{80^2(3)(27)^4}{108^4}\]

OpenStudy (hybrik):

o

OpenStudy (hybrik):

75

OpenStudy (freckles):

cool stuff

OpenStudy (freckles):

x+y+z=27+16+75

OpenStudy (hybrik):

42+75=127

OpenStudy (freckles):

so i guess they were integers after all

OpenStudy (hybrik):

What if they weren't what would you do?

OpenStudy (freckles):

isn't 27+16 equal to 43?

OpenStudy (freckles):

and isn't 43+75=118

OpenStudy (hybrik):

bad math confirmed.

OpenStudy (freckles):

well we just assumed they were real numbers in the beginning

OpenStudy (hybrik):

So thats why I got a 145/150 on this Mu Alpha contest

OpenStudy (freckles):

that includes the integers and non-integers

OpenStudy (freckles):

I was thinking that there should be a short cut for some reason but I do not see one

OpenStudy (hybrik):

Two awards, [Fan medal], https://www.youtube.com/watch?v=jH-miNEnIgE

OpenStudy (hybrik):

One more question maybe?

OpenStudy (freckles):

ok I will try to answer

OpenStudy (hybrik):

Given that a^2 sub 1 + a^2 sub 2 + a^2 sub 3= 7, find the maximum possible value of 3a sub 1 + 2a sub 2 + a sub 3

OpenStudy (freckles):

\[a^2_1+a^2_2+a^2_3=7 \\ \text{ find max of } 3a_1+2a_2+a_3\]

OpenStudy (hybrik):

yeah

OpenStudy (freckles):

I'm thinking that we can try to use that one thing it starts with an L lebniz?

OpenStudy (hybrik):

Leibniz, I am in 6th grade, I dont know this stuff well, explian more of it.

OpenStudy (freckles):

oh I was thinking of a calculus thing

OpenStudy (hybrik):

I do pre calculus

OpenStudy (freckles):

i'm so dumb it was another l word

OpenStudy (freckles):

lagrange multiplers

OpenStudy (hybrik):

What is that exactly?

OpenStudy (freckles):

http://tutorial.math.lamar.edu/Classes/CalcIII/LagrangeMultipliers.aspx this is what it is

OpenStudy (hybrik):

K then?

OpenStudy (freckles):

\[g(z,y,z)=x^2+y^2+z^2=7 \\ f(x,y,z)=3x+2y+z \\ f_x=3 \\ f_y=2 \\ f_z=1 \\ g_x=2x \\ g_y=2y \\ g_z=2z \\ \text{ so we have the equations } \\ 3= \lambda 2x \\ 2= \lambda 2y \\ 1 =\lambda 2z\]

OpenStudy (freckles):

we need to solve that bottom system for (x,y,z)

OpenStudy (freckles):

and we also have that first equation

OpenStudy (freckles):

\[3=\lambda 2x \\ 2=\lambda 2y \\ 1=\lambda 2z \\ x^2+y^2+z^2=7\]

OpenStudy (hybrik):

3x+2y+z = what number?

OpenStudy (freckles):

well if you can solve that system of equations above for (x,y,z) you can find out you might need to find lambda first though

OpenStudy (hybrik):

I honestly don't know what that is :(, I read it but it doesnt make my brain into order

OpenStudy (freckles):

so maybe we need another way to solve besides lagrange multiplers then

OpenStudy (freckles):

any ideas out there

OpenStudy (freckles):

that is the only one I have unfortunately

OpenStudy (freckles):

that lagrange multipler way is pretty cute I found lambda and it wasn't too bad it was just a quadratic equation and then as result I was able to find x,y, and z :(

OpenStudy (hybrik):

It is hard to learn pre cal in 6th grade :(

OpenStudy (freckles):

yeah I think lagrange multipler is a cal 3 method

OpenStudy (freckles):

or cal 2

OpenStudy (hybrik):

:_:

OpenStudy (hybrik):

I have, one more >_<

OpenStudy (hybrik):

Let x and y be real numbers satisfying 2/x = y/3 = x/y. Determine x^3

OpenStudy (hybrik):

Im assuming 0, but that just seems wrong

OpenStudy (freckles):

\[\frac{2}{x}=\frac{y}{3} \implies 6=xy \\ \frac{2}{x}=\frac{x}{y} \implies 2y=x^2 \\ \\ \frac{y}{3}=\frac{x}{y} \implies y^2=3x\] maybe we can play with these equations

imqwerty (imqwerty):

\[x^3+y^3+z^3=(x+y+z)^2-3(x^2y+y^2x+y^2z+z^2y+z^2x+x^2z)-9xyz\]

OpenStudy (hybrik):

@imqwerty the question i just asked is that the simplified part of it or the one before ">_<"

imqwerty (imqwerty):

is there any way to find\[x^3+y^3+z^3\]?

OpenStudy (hybrik):

@freckles Where did you get the values to equal from (6)

OpenStudy (freckles):

\[6=xy \\ 2y=x^2 \ \text{ note multiply both sides by } x \text{ and then use first equation }\]

imqwerty (imqwerty):

:O ok :)

OpenStudy (freckles):

\[\frac{2}{x}=\frac{y}{3} \\ \text{ multiply both sides by } 3 \\ \text{ multiply both sides by } x \\ \text{ this is how I got } 6=xy \] Is this the equation with 6 you are talking about @hybrik

OpenStudy (hybrik):

Or did you cross multiply 2 * 3 and x*y

OpenStudy (freckles):

that is what some people like to call it

OpenStudy (freckles):

that is what I did for each pair of fractions that were equal

OpenStudy (freckles):

I only needed the first two equations though

OpenStudy (freckles):

\[\frac{2}{x}=\frac{y}{3} \implies 6=xy \\ \frac{2}{x}=\frac{x}{y} \implies 2y=x^2 \\ \\ \frac{y}{3}=\frac{x}{y} \implies y^2=3x\ \] Do you see multiplying the second one on both sides by x gives 2xy=x^3 and you see that xy is given as 6 in the first equation?

OpenStudy (hybrik):

Im just going with my method, proceed your explanation

OpenStudy (freckles):

my method is done :p

OpenStudy (freckles):

you just replace xy with 6

OpenStudy (freckles):

2(6)=x^3